Consider the three points: A = (8,7) B = (1,5) C = (5,1). Determine the angle between AB and AC. =
Answer #1

solution Given poinls are A=1817) , B=(1,5) and c= (511) AB= (1-8,5-7)= (-7,-2) AC=(5-8, 1-7)=(-3,-6) the angle between AB ancoso = 2) +12 √53 545 § 33 O = COS (1905) 0 = 47.50